OpenStudy (anonymous):

They want me to write the problem in SIMPLEST RADICAL FORM. Am I doing it right?

OpenStudy (anonymous):

\[\sqrt{n}\sqrt[3]{n}\sqrt[6]{n}\] is this \[n ^{1/2}\]

OpenStudy (anonymous):

they are being multiplied

OpenStudy (anonymous):

no since \(\sqrt{n}=n^{\frac{1}{2}}\)

OpenStudy (anonymous):

Would the answer just be \[\sqrt{n}\sqrt[3]{n}\sqrt[6]{n}\] as the simplest it could go

OpenStudy (mathmale):

Hint: It may help you to see what needs to be done if you'd please re-write this expression as \[n ^{1/2}*n ^{1/3}*n ^{1/6}\]

OpenStudy (mathmale):

Hint: referring to "rules of exponents," what is\[x^a*x^b, ~when~ simplified?\]

OpenStudy (anonymous):

would it be \[n^{a+b}\] so 3/6 + 2/6 + 1/6 = 6/6, which is 1. So just n would be the answer?

OpenStudy (mathmale):

Yes, and that's very nice work on your part!

OpenStudy (anonymous):

Thanks, I appreciate the help! Could I ask you one more like this

OpenStudy (mathmale):

Sure, but would you please post it as a separate question.

OpenStudy (anonymous):

Ok! I'm about to, no problem